OpenStudy (anonymous):

Convert pi/20 radians into degrees.

OpenStudy (anonymous):

9

OpenStudy (anonymous):

There are 2 pi in a circle which is 360 degrees. pi/20 = 2pi/40 ---> 360/40 = 9

OpenStudy (anonymous):

What is the equation of the following graph? f(x) = cosθ f(x) = sinθ f(x) = tanθ f(x) = cotθ http://i729.photobucket.com/albums/ww292/akatushi101/broooo.gif

OpenStudy (anonymous):

tan

OpenStudy (anonymous):

Convert -245° into radians

OpenStudy (anonymous):

4.165

OpenStudy (anonymous):

soo what is the radian? like for example 49 pi/36

OpenStudy (anonymous):

use the relation ....1 radians = 57.2957795 degree 1 degree = 0.0174532925 radian

OpenStudy (anonymous):

or pi/180 = 1 radian

OpenStudy (anonymous):

and 180/pi=1 degree

OpenStudy (anonymous):

here is my choices 2pi/15.. -2pi/15 pi..-49pi/36 pi

OpenStudy (anonymous):

this is for which problem u giving me choices

OpenStudy (anonymous):

Convert -245° into radians

OpenStudy (anonymous):

oh...sorry its -49pi/36

OpenStudy (anonymous):

Given cos B = 11/8 find angle B in degrees. Round your answer to the nearest hundredth is it 52.33°?

OpenStudy (anonymous):

i guess it should be 57.31

OpenStudy (anonymous):

43.37° 52.33° 78.13° 65.55°

OpenStudy (anonymous):

those are the choicess

OpenStudy (anonymous):

ya then it is 52.33

OpenStudy (anonymous):

actually it looks like the question is wrong....because cos cannot exceed -1 and 1

OpenStudy (anonymous):

are u sure its 11/8 on the right side

OpenStudy (anonymous):

yes cos b

OpenStudy (anonymous):

ok i hope the answer is right and sorry if its wrong